OpenStudy (anonymous):

A sample of carbon-12 has a mass of 6.00 g. How many atoms of carbon-12 are in the sample? I have the answer but I don't understand how to solve it.

OpenStudy (anonymous):

Find the number of moles first by dividing the mass by the molar mass, then find the number of entities (atoms, molecules, electrons, etc.), by multiply the number of moles by Avogadro's constant, \[\text{entities} = \text{number of moles} \times (6.02 \times 10^{23})\] One mole of carbon-12 is exactly 12 g of carbon. 6 g of carbon-12 is half a mole of carbon-12. \[\text{atoms of C-12} = 0.500 ~\text{mol} \times (6.02 \times 10^{23})\]